OpenStudy (metalslayer):

A regular n-gon is inscribed in the unit circle. What is the perimeter for each n below? a.3 b.5 c.6 d.10 e.57 f.542 g. n h. The perimeter in part f should be close to what number? How close is it?

OpenStudy (anonymous):

Aye, mate. I may be of some assistance. But first, I need a favor from ye. Tell me when the observatory is.

OpenStudy (metalslayer):

What observatory?

OpenStudy (metalslayer):

I really need help on this

geerky42 (geerky42):

Do you know Trigonometry? Or at least Law of Cosines?

geerky42 (geerky42):

If so, this site may helps: http://mathcentral.uregina.ca/QQ/database/QQ.09.07/h/lindsay2.html

OpenStudy (metalslayer):

I've been on that website. I just don't get it. Can you please just solve for A and show the steps so I could do the rest? Thanks.

OpenStudy (metalslayer):

I do know trig and law of cosines and sines

geerky42 (geerky42):

What A?

OpenStudy (metalslayer):

a.3 b.5 c.6 d.10 e.57 f.542 g. n h. The perimeter in part f should be close to what number? How close is it?

OpenStudy (metalslayer):

Please show me how to answer some of them with the steps. Thank You.

OpenStudy (metalslayer):

By A i meant n=3

geerky42 (geerky42):

From site, our equation would be \(P = nc = n\sqrt{2-2\cos\left(\dfrac{360^\text o}{n}\right)}\)

geerky42 (geerky42):

Do you understand how this equation was derived?

OpenStudy (metalslayer):

P is perimeter, n is the n-gon, what is c?

geerky42 (geerky42):

c is side.

OpenStudy (metalslayer):

How do I find the side? I am only given the n.

geerky42 (geerky42):

You can use law of cosines to find the length of sides: